OpenStudy (anonymous):

3 square root 20 plus 2 square root five

OpenStudy (anonymous):

\[8\sqrt{5}\]

OpenStudy (anonymous):

oops sorry its \[8\sqrt{5}\]

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[3\sqrt{20}+2\sqrt{5}=3\sqrt{2*2*5}+2\sqrt{5}=6\sqrt{5}+2\sqrt{5}=8\sqrt{2}\]

OpenStudy (anonymous):

thank u :D

OpenStudy (anonymous):

what is the value of the expression-x(y-8)^2